OpenStudy (anonymous):

anti deriv. of 5/ (1+25s^2) ?

OpenStudy (anonymous):

5ln(1+25s^2)?

OpenStudy (anonymous):

no thats not it .. the derivative of that is 5arctan(25s)... i need the anti deriv. of 5/ (1+25s^2)

OpenStudy (anonymous):

That looks similar to the derivative of arctan. Which is du/(1+u2)=(arctanu)′ So know that, the u would be 5s so the answer would be arctan(5s). 5/(1+(5s)2)=(arctan5s)′

OpenStudy (anonymous):

correct me if i am wrong an antideriv is going from the derivative back to the function it's self?

OpenStudy (anonymous):

thanks scot.. haha i knew that .. yes thats right

OpenStudy (anonymous):

zbay, yes, its also called an integral.

OpenStudy (anonymous):

-5ln(u)? ,,,

OpenStudy (anonymous):

Anti-derivative of that?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

This, I think is one of the harder integrals to see how to do it. You have to use parts. Where u = lnx and dv = 1. Then du = 1/x and v = x. So uv - integral(vdu) is xlnx - x + C is the general form, and then times the -5 so -5xlnx + 5x + C is the final answer.

OpenStudy (anonymous):

wow..... ya i wasnt even close to that answer... but ya i get it now ... thanks